How do i Solve for X? x^(3/2)-26=99
Answer by mananth(16946)

x^(3/2)-26=99
x^(3/2)=99+26
x^(3/2)=125
Take the square
x^(3/2)^2 = (125)^2
x^3 = 15625
take the cube root
x= 25
...
m.ananth@hotmail.ca